OpenStudy (anonymous):

How long is the common chord of the circles x^2+y^2=4 and x^2+y^2=4x?

OpenStudy (anonymous):

\[x^2+y^2=4\] and \[x^2+y^2=4x\]

OpenStudy (anonymous):

The first circle has radius 2 and centered at the origin. The second one has radius 2 and centered at the (2,0)

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (anonymous):

How does \[x^2+y^2=4x\] become \[(x-2)^2+y^2=4?\]

OpenStudy (anonymous):

Because it seems like that's what you are implying

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry, but can you explain how?

OpenStudy (anonymous):

Expand the second equation

OpenStudy (anonymous):

I haven't expanded an equation at all since the summer, so I might be wrong but it seems like x^2+y^2=4x isn't expandable...

OpenStudy (anonymous):

\[ x^2 + y^2 =4 x\\ x^2 -4 x +y^2=0\\ x^2 -4 x + 4 - 4 + y^2 =0\\ (x-2)^2 +y^2 -4=0\\ (x-2)^2 +y^2 =4\\ \]

OpenStudy (anonymous):

This method is called: completing the square. Is it clear now

OpenStudy (anonymous):

Ahh it makes sense now, I should be able to do the rest. Thank you!

OpenStudy (anonymous):

YW